I cannot see the topics, sorry.
In my problem you have to use A=pi r^2.
Find the area of a circle if its diameter is 2 sqrt (84) centimeters. (two square roots of 84)
I have tried taking half of this and multiplying 2 of them together. I have tried multiplying the 2 times itself, and trying to then find the square root of it. I have the answer to the one before it, and I have tried everything I can to come out to the right answer. No matter how I do it, I cannot get it right.
I would like to know how to figure it out, if you would explain to me how to figure it out, I can find the answer. I just do not know how to even attempt to do it.

Answer by rfer(16322)

A=3.14*(2sqrt(84)/2)^2
A=3.14 (2*9.165/2)^2
A=3.14*84
A=263.76 sq cm